OpenStudy (anonymous):

A park has a triangular piece of land on the border.The altitude of the triangle is 500 feet.The base of the triangle is 240 feet. What is the area of this ptece of land?

OpenStudy (anonymous):

look up the formula for area of a triangle

OpenStudy (anonymous):

i did that i don't know

OpenStudy (campbell_st):

the altitude is also the perpendicular height needed to find the area here is a link to expain it. then \[Area = \frac{1}{2} \times base \times height\] just plug the values in and calculate the answer hope it helped

OpenStudy (campbell_st):

oops the link is http://www.mathwords.com/a/altitude_triangle.htm